You invest 10,000, part at 6% annual interest and the rest at 9% annual interest. If you received 684 in interest after one year, how much did you invest at each rate?
Answer by ptaylor(1894)

Let x=amount at at 6%
Then $10000-x=amount at 9%
Now we are told that the amount of interest received at 6% (.06x) plus the amount of interest received at 9% (.09(10000-x)) equals $684. So our equation to solve is:
.06x+.09(10000-x)=684 get rid of parens
.06x+900-.09x=684 subtract 900 from both sides
-.03x=-$216
x=$7200---------------------amount invested at 6%
10000-x=10000-7200=$2800-------------------amount invested at 9%

Ck
.06(7200)+.09(2800)=684
432+252=684
684=684

Hope this helps----ptaylor